OpenStudy (anonymous):

2r over 5r-5 = ? over 15r-15 PLEASE HELP

OpenStudy (anonymous):

6r

OpenStudy (anonymous):

can you show me how you got there? please

jhonyy9 (jhonyy9):

2r x x ----- = ------- = -------- ------ x = (2r(3(5r-5))/(5r-5)= 2r(3)=6r 5r-5 15r-15 3(5r-5) x=6r so hence 2r/(5r-5) = 6r/(15r-15)